please help me
im failing algebra!
Graph 2y=4x
m=____________(slope)
b=____________

Answer by Mathtut(3670)

I cannot get my graph to work
but the slope is found in the y=mx+b format where m is the slope and b is the y intercept
y=(4/2)x---->y=2x so slope is 2 and the y intercept is 0 . the graph is a line that goes thru the point (0,0) it extends at a 45 degree angle in the 1st quadrant
and also extends at a 45 degree angle in the 3rd quadrant...meaning it splits the 1st and 3rd quadrants through the point (0,0)
hope that helps